If a boat points its prow directly across a river, and its engine has one speed of 8 57 m/s What is its speed, relative to the river bank, if the river has a velocity of 3.22 mis? Number Units If the river is 52.9 m wide, how far down stream does it land when it reaches the opposite shore? Number Units
Answer #1


velocity of river relative to bank vrB = -3.22j

velocity of boat realtive to river Vbr = -8.57 i m/s

velocity of boat relative to bank VbB = vr + vb = -3.22j - 8.57 i


speed = sqrt(3.22^2 + 8.57^2) = 9.15 m/s


=============================


time taken to cross river t = 52.9/8.57


distance travelled down stream d = Vr*t = 3.22*52.9/8.57 = 19.87 m